A sample of sulfur dioxide occupies a volume of 561 mL at 48° C and 792 mm Hg. What volume will the sulfur dioxide occupy in ml at 4° C and 1 atm? Round to the nearest two decimal places.


1
Expert's answer
2021-02-25T03:44:03-0500

There are two stages of the gas:


At first stage

"V_1=561 ml"

"P_1=792 mmHg"

"T_1=48^o C= 321K"


At second stage:

"P_2=1atm=760mmHg"

"T_2=4^o C=277 K"

"V_2=V_2(let)"


Here we use Ideal gas eqn

PV=nRT

n="\dfrac{PV}{RT}"


In both the stages moles are constants, so


"\Rightarrow" "n_1=n_2"


"\Rightarrow" "\dfrac{P_1V_1}{RT_1}=\dfrac{P_2V_2}{RT_2}"

As R is constant both the sides,so


"\Rightarrow V_2=\dfrac{P_1V_1T_2}{T_1P_2}"


"\Rightarrow V_2=\dfrac{792\times561\times277}{321\times760}=" 504.49 ml